# InteLiPlan: An Interactive Lightweight LLM-Based Planner for Domestic Robot Autonomy

### v2

@@ Abstract
We introduce an interactive LLM-based framework designed to enhance the autonomy and robustness of domestic robots, targeting embodied intelligence. Our approach reduces reliance on large-scale data and incorporates a robot-agnostic pipeline that embodies an LLM. Our framework, InteLiPlan, ensures that the LLM’s decision-making capabilities are effectively aligned with robotic functions, enhancing operational robustness and adaptability, while our human-in-the-loop mechanism allows for real-time human intervention when user instruction is required. We evaluate our method in both simulation and on the real Toyota Human Support Robot (HSR). Our method achieves a 93% success rate in the ‘fetch me’ task completion with failure recovery, highlighting its capability in both failure reasoning and task planning. InteLiPlan achieves comparable performance to state-of-the-art large-scale LLM-based robotics planners, while using only real-time onboard computing. More information about InteLiPlan can be found at the project website: https://kimtienly.github.io/InteLiPlan.
@@ I Introduction
In recent years, the integration of artificial intelligence (AI) into robotics has led to significant advancements in automation and autonomous systems [1, 2, 3, 4, 5]. A key development in this field is the application of large language models (LLMs), such as GPT [6, 7], LLaMA [8, 9], and Mistral [10], which have proven to be powerful tools for understanding and generating human-like text. These models offer novel approaches for enhancing human-robot interaction and enabling more intuitive decision-making processes in robotics.
Despite their impressive capabilities, applying LLMs in robotics presents unique challenges [11]. These challenges stem primarily from the constraints of robot kinematics and the dynamic nature of the environments in which robots operate. Additionally, integrating these generative models within the existing robotics pipelines is non-trivial, as these are usually tightly coupled systems of perception, planning, and actuation. Having such systems running in real-time robotic deployment also poses a significant challenge due to the computational overhead. Moreover, robotic motion safety and robustness are critical concerns in human-robot environments. Therefore, developing a general-purpose autonomous robotic system for domestic scenarios remains a challenging open problem.
Figure 1: Step-by-step execution of InteLiPlan result on the physical HSR. In our system, the robot receives requests and guidance from humans, e.g., pick-and-place of an orange. Our lightweight onboard planner will generate the robot actions for the task as shown in the figure.
@@
An integral part of our framework is the human-in-the-loop (HITL) mechanism, which allows the system to handle system failures (e.g., partially observable environment, vision failures due to lightning, etc.) through human intervention. This feature is critical for maintaining autonomy, especially in complex environments where decisions by the LLM may fail to execute. The system allows human operators to input corrective actions directly, facilitating on-the-fly adaptation to new or unanticipated situations. This real-time interaction not only enhances the robustness and reliability of robotic operations but also enables fine-tuning of the LLM’s responses in accordance with human guidance.
@@
  • How can we achieve embodied AI with a lightweight planner that considers different components, including user commands, visual perception, and action feasibility?
  • How can robots make use of LLMs to translate human-language inputs to executable action directly?
@@ II-C Replanning for Failure Recovery
Failure recovery has been widely addressed in the literature for robotics planners. When it comes to LLM-based solutions, researchers typically address replanning from a second-level action check when the plan is generated [23, 24]. Vision-language models are also implemented to describe the error from the scene [25]. REFLECT [26] incorporates visual and audio sensory data to explain failures from the failed actions. This showcases the success of failure reasoning with a multimodal structure, however does not consider robotic constraints and remains a pure task planner. CAPE [27] instead proposes re-prompting to replan upon failure with preconditions similar to classical planning methods (e.g. PDDL, STRIPS). Considering the context of task and motion planning, this method works in the form of hierarchical layers and replans upon execution failure. In addition to recovering from execution, we designed a multimodal LLM-based planner that incorporates action feasibility during generating plans, which helps to reduce the time taken for re-querying the planner.
@@ III-A Problem Statement
Our work involves a mobile manipulator interacting with non-expert users to perform a human request. InteLiPlan shares the same set of skills with the motion planner, which includes one-step actions like go(), pick(), place(), open(), close(), search(), turn(). During operation, the robot receives input from the user $U$, which is a request for the first input, or guidance for recovery. The centralized planner, InteLiPlan, takes in the user input and the internal vision and motion verification functions, in order to generate a feasible plan. If a plan is found, the sequence of actions will be executed from a predefined skill library. If a failure is detected, the robot will notify the reason and ask for human instruction.
The robot obtains environmental information from a vision module, a state-of-the-art object detection model, processing input from the onboard RGB-D camera. The visual detection is represented as $O$. The motion verification module assesses the feasibility score $F$ of the action during planning.
The result is InteLiPlan, a real-time onboard LLM-based planner, with the ability to interact with humans in natural language and replan upon guidance. We evaluate InteLiPlan on domestic tasks, aiming at a human-centric and interpretable framework.
@@ III-B Interactive Planner with Multimodal Perception
Our planner uses a fine-tuned LLM model, which takes multimodal inputs from the user ($U$), visual observation ($O$), and motion feasibility ($F$), as shown in Fig. 2. Before feeding into the central planner, the system evaluates the inputs in the following order: $U$, $O$, $F$. Algorithm 1 explains the internal interaction between the modules. When the user sends a command to the central planner, the object list is extracted and sent to the vision module. The object detector runs as a service that detects objects in real-time, and stores a list of detected objects with their corresponding positions. When a task is sent to the robot, the system checks if the mentioned items are found in the current observation. If the object is found, the motion verification module gets the object position and checks if the target object is reachable. For example, in our implementation, we use the reachability graph designed in R-LGP[33], which is used as a feasilibity check service for LGP framework [34]. This reachability graph takes in the desired end-effector position (e.g. object location – $object\_loc$) and current end-effector position and determines whether or not there is a whole-body collision-free trajectory between the two end points with $F$.get_score($object\_loc$). In InteLiPlan, we incorporate a binary score for feasibility check, which applies, but not limited to the reachability graph in R-LGP. Last, we keep track of the conversation by inserting a history $H$ to each planner call. $H$ is initialized empty and records the human-robot interaction over time.
Algorithm 1 InteLiPlan pipeline
1:user $U$, robot $R$, interaction history $H$,
2:        vision module $O$, feasibility module $F$
3:inputs $\leftarrow U$.input()
4:objects_list $\leftarrow$ $R$.extract(inputs)
5:H $\leftarrow$ None
6:while not ($R$.task_is_complete() or time_out()) do
7:     objects_loc $\leftarrow$ $O$.get_position(objects_list)
8:     result $\leftarrow$ $R$.plan($H$, inputs, $O$.is_seen(objects_list),
9:                          $F$.get_score(objects_loc))
10:     if $R$.get_failure(resultthen
11:         $H\leftarrow H$ + inputs + $R$.get_failure(result)
12:         inputs $\leftarrow U$.get_guidance()
13:     else
14:         $R$.execute($R$.get_plan(result))
InteLiPlan is formulated as $A=\phi(H,U,O,F)$, where $A=\{a_{t},...,a_{t+T}\}$ represents the output sequence of actions. While $U$ is the direct input from the user, $H$, $O$ and $F$ are the internal conversations of the system. $H$ is a quoted string that tracks the previous conversation for contextual information. We design $O$ as a text-based feedback from the visual observation, and $F\in\{0,1\}$ as a binary signal indicating the feasibility of the action. The multimodal inputs to the planner are represented in specific tokens: <history>, <user>, <vision>, <feasibility>, where the token for the output is <robot>. One example of the multimodal conversation is shown in Fig. 1.
Our framework features a centralized LLM-based planner to utilize the integration and coordination between submodules. We provide a plug-and-play modular design to leverage state-of-the-art models across various research domains (e.g. object recognition) in a zero-shot way.
@@ III-C Failure Recovery
We consider three main categories of failures: planning failure, task confusion and execution failure, as outlined in Table I. The multimodal structure allows the planner to reason around failures through feedback from the vision and motion verification modules, extracting the missing information to complete the task. When a failure occurs, our system can report to the human with a contextual description, interpreting the current observation and failure that occurred. By fine-tuning the LLM, the model can intelligently process human inputs and replan based on the human guidance. For instance, vision failures due to partial observation or adverse lighting conditions can be mitigated with human insights, leveraging the planner’s capability to maintain a record of initial commands through the internal dialogue $H$ and to replan upon system failure as necessary, as demonstrated in Algorithm 1. This capability significantly enhances transparency, robustness and reliability. By injecting fine-tuning data, the output is guaranteed to consider necessary observations before executing any actions.
@@ III-D Text-only Fine-Tuning Data for Robotics Planner
To fine-tune the LLM model, we collected a pure-text customized dataset that includes scenarios both with and without failure recovery, where the interactions in failure cases are limited to a maximum of two rounds. The dataset is formulated as $D=\{H_{i},U_{i},O_{i},F_{i},R_{i}|i=0,1,2,...N\}$. We loop through a list of objects and commands to generate the command and expected sequence of actions. Visual observation $O_{i}$ and feasibility score $F_{i}$ is provided in the form of text. This text-based-only fine-tuning approach enables the model to easily adapt to different robots without further modifications or re-parameterization.
Our methodology employs Low-Rank Adaptation (LoRA) [36] to fine-tune LLM efficiently. Instead of updating all model parameters, LoRA introduces trainable low-rank matrices into key projection layers, significantly reducing computational and memory overhead. LoRA is one of the PEFT [32] techniques that has also been proven to be superior over full fine-tuning as a parameter-efficient fine-tuning method in [37]. We implement supervised fine-tuning with the collected dataset, where $\{H_{i},U_{i},O_{i},F_{i}\}$ is the input, and $R_{i}$ is the desired output. This approach enables efficient adaptation of the model to domain-specific tasks while maintaining scalability.
@@ IV Experimental Results
Experimental Setup. We evaluate our approach using the Toyota Human Support Robot (HSR) in a domestic environment. The first set of experiments is a ‘fetch me’ task, in an environment that includes 10 seen objects and 20 unseen objects, and the planner is evaluated on 4 seen request commands, 23 unseen request commands, and replan with 9 seen guidance commands, and 18 unseen guidance commands. The unseen commands represent the variants in human-like conversation. We then scale up the model to test with the 101 task requests from SayCan [1]. The vision module receives inputs from the RGB-D camera of the HSR, then processes object recognition with YOLO [38] and uses pose estimation to obtain object locations. In this way, the robot can automatically verify the presence and precise location of objects relative to the robot’s map frame. For feasibility verification, we use the reachability graph from [33], a sampling-based approach that efficiently checks for paths from a starting point to a goal while ensuring feasible configurations for the robot reaching towards the target object. The LLM model for our approach is fine-tuned from Mistral 7B [10] and Deepseek 8B [39], state-of-the-art lightweight LLM models, which is suitable for onboard deployment. The fine-tuning data includes all failure categories outlined in Table I. Before the execution, the robot is provided with a map of the room, e.g. large furniture like tables or drawers, and a list of motion APIs that are shared between the planner and the controller.
@@
Metrics. To assess the performance of the proposed method we use five metrics. 1) Task planning success rate: the ratio of the no-failure trials where the robot successfully generates the action sequence by understanding the assigned task. 2) Failure explanation success rate: the ratio of the trials where the robot successfully generates the reason for not being able to find the action sequence. 3) Failure recovery success rate: the ratio of the trials where the robot successfully generates the recovery plan in response to human guidance. 4) Execution success rate: the ratio of the trials where the robot successfully executes the given request. 5) Inference time: the time taken to proceed the formulated multimodal inputs, showcasing the real-time capability of the model.
Baselines. We compare Mistral-Ours and Deepseek-Ours with few-shot prompting on Mistral - Mistral-Prompt and Deepseek - Deepseek-Prompt. The samples are in the same form as the dataset for fine-tuning, as discussed in Section III-D. The second comparison is with SayCan [1]. Given that this method does not have failure recovery ability, we only report the result during the scalability experiments. Generative result from PaLM 8B [40] is also reported as a lightweight LLM baseline.
IV-A Comparison of Different Methods in Feasible Tasks
@@
In our first experiment, we assume successful outcomes in both object detection and robotic motion feasibility. Here, the planner is solely required to generate the appropriate action sequences based on human inputs. An example of such conversation is provided in Fig. 3a.
@@
Despite being told to list a sequence of actions only, Mistral-Prompt and Deepseek-Prompt outputs include redundant text. For a fairer comparison, we only extracted the outputted action sequence when determining the success rate, and ignored the lengthy explanation and slightly misused words (e.g. ‘pick up’ vs ‘pick’) for action API interface. As presented in the w/o failure column of Table II, the results demonstrate a superior performance of our models, achieving 100% success in scenarios with seen commands and objects, and maintaining high effectiveness even with unseen objects or commands. In contrast, Mistral-Prompt and Deepseek-Prompt experienced performance drops in scenarios involving unseen elements, with more sensitivity over unseen textual commands. It is also observed that Deepseek-Prompt behaves better than Mistral-Prompt, while Deepseek-Ours has slightly lower success rates than Mistral-Prompt.
IV-B Evaluation of Robustness in Failure Explanation
@@
In this test, we break down the evaluation of failure explanation in the failure recovery process, investigating the ability of the model to explicitly describe what hinders the process of planning actions. The failure experiments include all categories outlined in Table I, with examples of ‘planning failure’ detection in Fig. 3b,c. We only include ‘execution failure’ in failure recovery experiments (Section IV-C), as it is a failure description that is triggered from the controller level.
@@
IV-C Evaluation of Robustness in Failure Recovery

### v3

@@ Abstract
We introduce an interactive LLM-based framework designed to enhance the autonomy and robustness of domestic robots, targeting embodied intelligence. Our approach reduces reliance on large-scale data and incorporates a robot-agnostic pipeline that embodies an LLM. Our framework, InteLiPlan, ensures that the LLM’s decision-making capabilities are effectively aligned with robotic functions, enhancing operational robustness and adaptability, while our human-in-the-loop mechanism allows for real-time human intervention when user instruction is required. We evaluate our method in both simulation and on the real Toyota Human Support Robot and Anymal D-Unitree Z1 platforms. Our method achieves a 95% success rate in the ‘fetch me’ task completion with failure recovery, highlighting its capability in both failure reasoning and task planning. InteLiPlan achieves comparable performance to state-of-the-art large-scale LLM-based robotics planners, while using only real-time onboard computing.
@@ I Introduction
In recent years, incorporating large language models (LLMs) and vision-language models (VLMs) into robotic systems as high-level planners has made remarkable progress in the embodied AI field, with applications in domestic autonomy and robotic assistance. In such hierarchical frameworks, multimodal perception modules are integrated into the system, and LLMs or VLMs serve as central components for scene understanding and reasoning, as well as generating action steps. Exploiting the reasoning abilities of these advanced models, LLM-based frameworks have shown promise in enhancing human-robot interaction and enabling intuitive decision-making in robotics.
Despite their impressive capabilities, applying LLMs in robotics presents unique challenges [1]. These challenges stem primarily from the constraints of robot kinematics and the dynamic nature of the environments in which robots operate. Additionally, integrating these generative models within the existing robotics pipelines is non-trivial, as these are usually tightly coupled systems of perception, planning, and actuation. Having such systems running in real-time robotic deployment also poses a significant challenge due to the computational overhead. Moreover, robotic motion safety and robustness are critical concerns in human-robot environments. Therefore, developing a general-purpose autonomous robotic system for domestic scenarios remains a challenging and open problem.
Figure 1: Step-by-step execution of InteLiPlan result on the physical Anymal D-Unitree Z1. In our system, the robot receives either requests or guidance from humans. Our lightweight onboard planner will generate the robot actions and execute them sequentially.
@@
An integral part of our framework is the human-in-the-loop (HITL) mechanism, which allows the system to handle system failures (e.g., partially observable environment, vision failures due to lightning, etc.) or task confusions through human intervention. This feature is critical for maintaining autonomy, especially in complex environments where decisions by the LLM may fail to reason what the user requested from an embodied perspective. The system allows human operators to input corrective actions directly, facilitating on-the-fly adaptation to unanticipated situations. This real-time interaction not only enhances the robustness and reliability of robotic operations but also enables adjusting the LLM’s responses in accordance with human guidance.
@@
  • How can we design a lightweight LLM-based planner that effectively integrates user commands, visual perception, and action feasibility with onboard computing resources?
  • How can robots make use of LLMs to translate human-language inputs to reliable and executable action directly?
@@
  • How can reachability-aware LLM-based robotic planners be applicable cross-embodiment?
@@ II-C Replanning for Failure Recovery
Failure recovery has been widely addressed in the literature for robotics planners. When it comes to LLM-based solutions, researchers typically address replanning from a second-level action check when the plan is generated [19, 20]. Vision-language models are also implemented to describe the error from the scene [21]. REFLECT [22] incorporates visual and audio sensory data to explain failures from the failed actions. This showcases the success of failure reasoning with a multimodal structure, however does not consider robotic constraints and remains a pure task planner. ReplanVLM [23] uses three different VLM agents for planning, replanning with internal and external errors, indicating that robust solution is hard to obtain by prompting one single LLM planner. CAPE [24] instead proposes re-prompting to replan upon failure with preconditions similar to classical planning methods (e.g. PDDL, STRIPS). Considering the context of task and motion planning, this method works in the form of hierarchical layers and replans upon execution failure. In addition to recovering from execution, we designed a multimodal LLM-based planner that incorporates action feasibility during generating plans, which helps to reduce the time taken for re-querying the planner.
@@ III-A Problem Statement
[…]
The robot obtains environmental information from a vision module, a state-of-the-art object detection model, processing input from the onboard RGB-D camera. The visual detection is represented as $O$. The motion verification module assesses the feasibility score of the action during planning. In this setting, we define InteLiPlan as a real-time onboard LLM-based planner, with the ability to interact with humans in natural language and replan upon guidance. We evaluate InteLiPlan on domestic tasks, aiming at a human-centric and interpretable framework.
@@ III-B Interactive Planner with Multimodal Perception
Our system processes multimodal inputs from the user $T_{user}$, visual observations $O$, and motion feasibility scores $f_{score}$. At the core of our system is a fine-tuned LLM that serves as the interactive planner, as illustrated in Fig. 2. Algorithm in Fig. 2 explains the internal interaction between the modules. When the user sends a command to the central planner, the request is first sent to the vision module. The vision module runs as a service that detects objects in real-time, and continuously update a list of detected objects with their corresponding positions in the vision memory. When a task is sent to the robot, the system checks if the mentioned items are found in the current observation. If the object is found, the motion verification module gets the object position and checks if the target object is reachable. For example, in our implementation, we use the reachability graph designed in R-LGP[30], which is used as a feasilibity check service for logic-geometric programming (LGP) framework [31]. This reachability graph takes in the desired end-effector position (e.g. object location $x_{obj}$) and current end-effector position and determines whether or not there is a whole-body collision-free trajectory between the two end points, regarded as $f_{score}(x_{obj},G_{map})$. In InteLiPlan, we incorporate a binary score for feasibility check, which applies, but not limited to the reachability graph in R-LGP. Lastly, we keep track of the conversation by inserting a history $T_{hist}$ to each planner call. $T_{hist}$ is initialized empty and records the initial request with the most recent failure detected by the planner.
Bringing together multimodal observations and interaction history, we formally define InteLiPlan as:
$A=\phi(T_{hist},T_{user},f_{seen}(l_{obj}),f_{score}(x_{obj},G_{map})),$ (1)
[…]
Our framework features a centralized LLM-based planner to utilize the integration and coordination between submodules. We provide a plug-and-play modular design to leverage state-of-the-art models across various research domains (e.g., object recognition) in a zero-shot way.
@@ III-C Failure Recovery
We consider three main categories of failures: planning failure, task confusion and execution failure, as outlined in Table I. The multimodal structure allows the planner to reason around failures through feedback from the vision and motion verification modules, extracting the missing information to complete the task. When a failure occurs, our system can report to the human with a contextual description, interpreting the current observation and failure that occurred. By fine-tuning the LLM, the model can intelligently process human inputs and replan based on human guidance. For instance, vision failures due to partial observation or adverse lighting conditions can be mitigated with human insights, leveraging the planner’s capability to maintain a record of initial commands through the internal dialogue $T_{hist}$ and to replan upon system failure as necessary. This capability significantly enhances transparency, robustness, and reliability. By injecting fine-tuning data, the output is guaranteed to consider necessary observations before executing actions.
@@ III-D Text-only Fine-Tuning Data for Robotics Planner
To fine-tune the LLM model, we collected a pure-text customized dataset that includes scenarios both with and without failure recovery, where the interactions in failure cases are limited to a maximum of two rounds. The dataset is formulated as $D=\{T_{hist},T_{user},f_{seen},f_{score},A\}$. We loop through a list of objects and commands to generate the command and expected sequence of actions. Visual observation $f_{seen}$ and feasibility score $f_{score}$ will be processed by the fixed visual module and feasibility module to generate textual signals. This text-based-only fine-tuning approach enables the model to easily adapt to different robots without further modifications or re-parameterization.
Our methodology employs Low-Rank Adaptation (LoRA) [33] to fine-tune LLM efficiently. Instead of updating all model parameters, LoRA introduces trainable low-rank matrices into key projection layers, significantly reducing computational and memory overhead. LoRA is one of the PEFT [29] techniques that has also been proven to be superior over full fine-tuning as a parameter-efficient fine-tuning method in [34]. We implement supervised fine-tuning with the collected dataset, where $\{T_{hist},T_{user},f_{seen},f_{score}\}$ is the input, and $A$ is the desired output, with cross-entropy loss. This approach enables efficient adaptation of the model to domain-specific tasks while maintaining scalability.
@@ IV-A Evaluation of Robustness Across Different Failures
In the first set of experiment, we evaluate our approach using the Toyota Human Support Robot (HSR) in a ‘fetch me’ task. The environmental setup includes 10 seen objects and 20 unseen objects, and the planner is evaluated on 4 seen request commands, 23 unseen request commands, and replan with 9 seen guidance commands, and 18 unseen guidance commands. The unseen commands represent the variants in human-like conversation. The vision module receives inputs from the RGB-D camera of the HSR, then processes object recognition with YOLO [35] and uses pose estimation to obtain object locations. For feasibility verification, we use the reachability graph from [30], a sampling-based approach that efficiently checks for paths from a starting point to a goal while ensuring feasible configurations for the robot reaching towards the target object. To verify applicability, the LLM model for our approach is fine-tuned from Mistral 7B [36] and Deepseek 8B [37], state-of-the-art lightweight LLM models, which is suitable for onboard deployment. The fine-tuning data includes 300 samples covering all failure categories outlined in Table I. Before the execution, the robot is provided with a map of the room, e.g. large furniture like tables or drawers, and a list of motion APIs that are shared between the planner and the controller.
@@
To assess the performance of the proposed method in this experiment, we use three metrics. 1) Task planning success rate: the ratio of the no-failure trials where the robot successfully generates the action sequence by understanding the assigned task. 2) Failure explanation success rate: the ratio of the trials where the robot successfully generates the reason for not being able to find the action sequence. 3) Failure recovery success rate: the ratio of the trials where the robot successfully generates the recovery plan in response to human guidance.
We compare Mistral-Ours and Deepseek-Ours with few-shot prompting on Mistral - Mistral-Prompt and Deepseek - Deepseek-Prompt. The samples are in the same form as the dataset for fine-tuning, as discussed in Section III-D.
IV-A1 Comparison of Different Methods in Feasible Tasks
@@
In our first experiment, we assume successful outcomes in both object detection and robotic motion feasibility. Here, the planner is solely required to generate the appropriate action sequences based on human inputs. An example of such conversations is provided in Fig. 3a.
@@
Despite explicit instructions to produce solely an action sequence, Mistral-Prompt and Deepseek-Prompt often generate extraneous text. For a fairer comparison, we only extracted the outputted action sequence when determining the success rate, and ignored the lengthy explanation and slightly misused words (e.g., ‘pick up’ vs ‘pick’) for the action API interface. As presented in the w/o failure column of Table II, the results demonstrate a superior performance of our models, achieving 100% success in scenarios with seen commands and objects, and maintaining high effectiveness even with unseen objects or commands. In contrast, Mistral-Prompt and Deepseek-Prompt experienced performance drops in scenarios involving unseen elements, with more sensitivity over unseen textual commands. It is also observed that Deepseek-Prompt behaves better than Mistral-Prompt, while Deepseek-Ours has slightly lower success rates than Mistral-Ours.
IV-A2 Evaluation of Robustness in Failure Explanation
@@
In this test, we break down the evaluation of failure explanation in the failure recovery process, investigating the ability of the model to explicitly describe what hinders the process of planning actions. The failure experiments include all categories outlined in Table I, with examples of ‘planning failure’ detection in Fig. 3b,c. We only include ‘execution failure’ in failure recovery experiments (Section IV-A3), as it is a failure description that is triggered from the controller level.
@@
IV-A3 Evaluation of Robustness in Failure Recovery